# ASH: Animatable Gaussian Splats for Efficient and Photoreal Human Rendering

### v1

@@ Appendix A Overview
In this appendix, we provide more details regarding the following aspects: More implementation details (Sec. B); more ablative studies (Sec. C); runtime analysis for the major components (Sec. D); realtime applications built upon ASH (Sec. E); more detailed discussion on limitations and future directions (Sec. F).
@@ Appendix B Implementation Details
Deformable Template Mesh. We adopt the formulation mentioned in Habermann et al. [13] for modeling the deformable template mesh, which deforms the template mesh vertices $\mathbf{V}_{\mathrm{m}}$ in the canonical space with a learned embedded deformation [55, 52]:
@@
where $\bar{\mathbf{V}}_{f,i}\in\mathbb{R}^{3}$ denotes the deformed template vertices in the rest pose. $\mathcal{N}_{\mathrm{nv},i}\in\mathbb{N}$ indicates the indexes for the embedded graph node that are connected to the $i$ th vertex on the template mesh. $\mathbf{V}_{\mathrm{G},j}\in\mathbb{R}^{3}$, $\mathbf{A}_{j}\in\mathbb{R}^{3}$, and $\mathbf{T}_{j}\in\mathbb{R}^{3}$ stands for the rest positions, rotation Euler angles, and translations of the embedded graph nodes. Notably, the connectivity of the embedded graph $\mathbf{V}_{\mathrm{G},j}$ can be adopted by simplifying the template mesh $M$ using quadric edge collapse decimation [8, 12]. Moreover, the connection, as well as the connection weights $w_{i,j}$, between the template mesh $\mathbf{V}_{\mathrm{m}}$ and the embedded graph are generated following the configuration in Sumner et al. [55]. $R(\cdot)\in\mathbb{R}^{3\times 3}$ denotes the function that converts the Euler angle to a rotation matrix. $\mathbf{D}_{i}\in\mathbb{R}^{3}$ indicates the per-vertex displacement to model an even finer level of geometry details. Specifically, embedded graph parameters $\mathbf{V}_{\mathrm{G},j}$, $\mathbf{A}_{j}$, and per-vertex displacements $\mathbf{D}_{i}$ are derived from skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ with structure-aware graph convolution neural networks. We refer to Harbermann et al. [13] for more details.
@@ Appendix C Ablations
Number of Camera Views. To assess the robustness of ASH against sparser camera view supervision, we conducted ablative experiments that take multi-view videos from $12$, $30$, and $60$ cameras as supervision, termed as w/ 12.cam, w/ 30.cam, and w/ 60.cam. Note that the selected camera views are evenly distributed in the studio. As illustrated in Fig. 1 and Tab. 1, ASH can still accurately synthesize the imagery of the animatable characters when training with sparser input views.
The Impact of 2D Learning. To validate the efficacy of the 2D texel paradigm for 3D Gaussian splats, we conducted an ablative experiment that predicts the 3D Gaussian parameters directly from 3D, termed as w/ MLP. We adopted an 8-layer MLP that consumes the skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ and the positional-encoded canonical Gaussian position $\boldsymbol{\bar{\mu}}_{i}$, predicting the Gaussian splat parameters $\{\mathcal{G}_{i}\}$ in the canonical space. Specifically, the width for each hidden layer of the MLP is set to $256$. Similar to ASH, the canonical Gaussian splats are wrapped to observation space through DQ skinning [21]. As illustrated in Fig. 1 and Tab. 1, directly learning the Gaussian parameters in 3D will lead to blurry rendering and cannot preserve the motion-dependent wrinkle details. In contrast, ASH, which formulates the learning of 3D Gaussian splats as image translation in 2D texel space, delivers high-quality rendering with delicate details.
@@
Appendix D Runtime Analysis
@@
In this section, we conduct a detailed runtime analysis for each major component in ASH. Specifically, we record the runtime for each major component when rendering a 1K ($1285\times 940$) image on a single Nvidia Tesla A100 graphics device. Additionally, the runtime analysis is benchmarked on models with different texture space resolution, specifically at $128$, $512$, and $256$, referred to as w/ 128.res, w/ 512.res, and Ours, respectively. Here, we divide the rendering pipeline of ASH into four steps as follows:
@@
  • Deriving motion-aware texture maps $(\mathbf{T}_{\mathrm{n},f},\mathbf{T}_{\mathrm{p},f})$ from deformable template meshes $M(\boldsymbol{\theta}_{f})$, termed as Stg.2.
@@ Appendix E Application
Fig. 2 presents a screenshot of ASH Player, which runs in the web browser on a personal computer. The backend model of ASH Player, i.e., ASH, is deployed on the GPU cluster server. Once users specify the skeletal poses and virtual camera views, ASH Player will present the photoreal rendering of animatable characters, which is real-time computed and streamed from the GPU cluster server. Moreover, ASH Player allows users to inspect the animatable characters with spiral camera views. Please refer to the supplementary video for a more comprehensive visualization of ASH Player.
@@
Appendix F Limitations
Although ASH enables the high-fidelity, real-time rendering of animatable human characters, it has certain limitations that we hope to address in the future.
Firstly, ASH does not extract detailed explicit geometry from the Gaussian splats. We will explore refining the explicit template meshes by backpropagating the gradient from image space into the template meshes using splatting. Additionally, ASH does not model topological changes like opening a jacket. Future research might focus on modeling the topological changes with the adaptive adding and removal of Gaussian splats introduced in the original 3D Gaussian splatting paper [22].

### v2

@@ Appendix A Overview
In this appendix, we provide more details regarding the following aspects: More implementation details (Sec. B); more ablative studies (Sec. C); more results with driven poses from novel datasets (Sec. D); runtime analysis for the major components (Sec. E); realtime applications built upon ASH (Sec. F); more detailed discussion on limitations and future directions (Sec. G).
@@ Appendix B Implementation Details
Deformable Template Mesh. We adopt the formulation introduced in Habermann et al. [15] for modeling the deformable template mesh, which deforms the template mesh vertices $\mathbf{V}_{\mathrm{m}}$ in the canonical space with a learned embedded deformation [63, 60]:
@@
where $\bar{\mathbf{V}}_{f,i}\in\mathbb{R}^{3}$ denotes the deformed template vertices in the rest pose. $\mathcal{N}_{\mathrm{nv},i}\in\mathbb{N}$ indicates the indices for the embedded graph node [63] that are connected to the $i$ th vertex on the template mesh. $\mathbf{V}_{\mathrm{G},j}\in\mathbb{R}^{3}$, $\mathbf{A}_{j}\in\mathbb{R}^{3}$, and $\mathbf{T}_{j}\in\mathbb{R}^{3}$ denotes the rest positions, Euler angles, and translations of the embedded graph nodes. Notably, the connectivity of the embedded graph $\mathbf{V}_{\mathrm{G},j}$ can be adopted by simplifying the template mesh $M$ using quadric edge collapse decimation [8, 12]. Moreover, the connection, as well as the connection weights $w_{i,j}$, between the template mesh $\mathbf{V}_{\mathrm{m}}$ and the embedded graph are generated Sumner et al. [63]. $R(\cdot)\in\mathbb{R}^{3\times 3}$ denotes the function that converts the Euler angle to a rotation matrix. $\mathbf{D}_{i}\in\mathbb{R}^{3}$ indicates the per-vertex displacement to model an even finer level of geometry details. Specifically, embedded graph parameters $\mathbf{V}_{\mathrm{G},j}$, $\mathbf{A}_{j}$, and per-vertex displacements $\mathbf{D}_{i}$ are derived from skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ with structure-aware graph convolution neural networks. We refer to Harbermann et al. [15] for more details.
@@ Appendix C Ablations
Number of Camera Views. To assess the robustness of ASH against sparser camera view supervision, we conducted ablative experiments that take multi-view videos from $12$, $30$, and $60$ cameras as supervision, termed as w/ 12.cam, w/ 30.cam, and w/ 60.cam. Note that the selected camera views are evenly distributed in the studio. As illustrated in Fig. 1 and Tab. 1, ASH can still accurately synthesize the animatable characters when training with sparser input views.
The Impact of 2D Learning. To validate the efficacy of the 2D texel paradigm for 3D Gaussian splats, we conducted an ablative experiment that predicts the 3D Gaussian parameters directly from 3D, termed as w/ MLP. We adopted an 8-layer MLP that consumes the skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ and the positional-encoded canonical Gaussian position $\boldsymbol{\bar{\mu}}_{i}$, predicting the Gaussian splat parameters $\{\mathcal{G}_{i}\}$ in the canonical space. Specifically, the width for the hidden layers of the MLP is set to $256$. Similar to ASH, the canonical Gaussian splats are transformed to observation space through Dual Quaternion skinning [25]. As illustrated in Fig. 1 and Tab. 1, directly learning the Gaussian parameters in 3D will lead to blurry rendering and cannot preserve the motion-dependent wrinkle details. In contrast, ASH, which formulates the learning of 3D Gaussian splats as image translation in 2D texel space, delivers high-quality rendering with delicate details.
@@
The Reliance on the Accuracy of the Template. Although our method is conditioned on a template mesh, it can compensate for tracking errors with learnable motion-aware residual deformations for the Gaussian splats. To validate our method’s robustness against errors in mesh tracking, we replaced the original template mesh with SMPL body meshes [38]. Despite large deviations between the template and "the real surface," our method generates visually plausible results (Fig. 2) and achieves significantly better quantitative performance than the SOTA real-time methods (Tab. 2), which heavily relies on an accurate deformable template [15].
Training Pose Testing Pose
Methods PSNR LPIPS PSNR LPIPS
SMPL 34.93 14.90 28.75 24.87
Ours 35.47 8.30 27.13 20.22
Table 2: ASH conditioned on SMPL. ASH achieves significantly better quantitative performance than the SOTA real-time methods.
Figure 2: ASH conditioned on SMPL. Despite large deviations between the underlying template and the real surface, ASH generates visually plausible results.
Appendix D More Results
In Tab. 2 in the main paper, we report the quantitative and qualitative performance on the testing set of the DynaCap dataset, which is an established and challenging benchmark, with the testing set containing more than $7000$ frames showing strongly varying poses.
To further highlight the pose generalization ability of ASH, we retarget our skeleton to SMPL motions from the AMASS dataset (DanceDB) [40] to drive our character. Fig. 3 illustrates that even for motions from an entirely different dataset, ASH could generate photoreal rendering with delicated wrinkle details.
Figure 3: Results with AMASS DanceDB motion. ASH produces photorealistic rendering given the motion from an entirely different dataset.
Appendix E Runtime Analysis
@@
In this section, we conduct a detailed runtime analysis for each major component in ASH. Specifically, we record the runtime for each major component when rendering a 1K ($1285\times 940$) image on a single Nvidia Tesla A100 graphics device. Additionally, the runtime analysis is benchmarked on models with different texture space resolution, specifically at $128$, $512$, and $256$, referred to as w/ 128.res, w/ 512.res, and Ours, respectively. Here, we divide the rendering pipeline of ASH into four steps:
@@
  • Computing motion-aware texture maps $(\mathbf{T}_{\mathrm{n},f},\mathbf{T}_{\mathrm{p},f})$ from deformable template meshes $M(\boldsymbol{\theta}_{f})$, termed as Stg.2.
@@ Appendix F Application
Fig. 4 presents a screenshot of ASH Player, which runs in the web browser on a personal computer. The backend model of ASH Player, i.e., ASH, is deployed on the GPU cluster server. Once users specify the skeletal poses and virtual camera views, ASH Player will present the photoreal rendering of animatable characters, which is real-time computed and streamed from the GPU cluster server. Moreover, ASH Player allows users to inspect the animatable characters with spiral camera views. Please refer to the supplementary video for a more comprehensive visualization.
@@
Appendix G Limitations
Although ASH enables high-fidelity, real-time rendering of animatable human characters, it has certain limitations that we hope to address in the future. Firstly, ASH does not extract detailed explicit geometry from the Gaussian splats. We will explore refining the explicit template meshes by backpropagating the gradient from image space into the template meshes using splatting. Additionally, ASH does not model topological changes like opening a jacket. Future research might focus on modeling the topological changes with the adaptive adding and removal of Gaussian splats introduced in the original 3D Gaussian splatting paper [26]. Lastly, as various factors could affect the appearance of dynamic clothed humans, it is unfeasible to establish a one-to-one correspondence between the skeletal motions and the dynamic clothed human appearance. Future research will explore different types of fine-grained control to define human rendering, e.g., the external physical forces.